OpenStudy (anonymous):

(x - 3) - 2(x + 6) = -5

OpenStudy (snuggielad):

First you need to use the distributing...do you know how to do that? lets not give answers moli...

OpenStudy (anonymous):

GO TO THIS WEB SITE IT GIVE U THE ANSWERS MATHWAY.COM

OpenStudy (snuggielad):

again...both of you...welcome to openstudy but please do not send them to sites where they can cheat either this is a cheat free zone

OpenStudy (anonymous):

oops sorry :p

OpenStudy (anonymous):

distribute the 2 to get -2x+12 then add/subtract the x-3 to get -x+9 subtract 9 from each side you get -x=-4 so x=4

OpenStudy (anonymous):

then to check you can plug it in (4-3)-2(x+6) and see if it equals -5

OpenStudy (anonymous):

2(4+6) my bad

OpenStudy (anonymous):

and it doesnt so i messed up. ignore me! lol

OpenStudy (anonymous):

ahhan. . first u gonna solve -2(x+6) which is -2x-12 now solve the LHS x-3-2x-12 which comes as -x-15 so -x-15=-5 by transposition -x=15-5 so x=-10 :p

OpenStudy (anonymous):

it's ok @mathgeek101